# Block-Attention for Efficient RAG

### v4

@@ 3.5 Main Results
From the results in Table 1, we can draw three important conclusions:
1) It is not advisable to directly switch from self-attention to Block-Attention, as it will lead to a sharp accuracy drop. For example, removing the Block fine-tune process leads to the 21.99% average absolute performance decrease on all four benchmarks for Llama3-8B and Mistral-7B models.
2) However, if we use the Block-Attention mechanism in the fine-tuning stage, then the resultant model has almost the same performance as the self-attention model or is even slightly better on some datasets. For instance, Mistral-block-ft outperforms the strong baseline Mistral-vanilla-sft, with 2.7% absolute improvements on four benchmarks.
@@
3) The position re-encoding operations are essential for the Block-Attention Model. Removing it leads to significant performance drop—average 4% decreases in accuracy on four datasets.
Figure 4: The accuracy of model checkpoint at different steps
The first conclusion is easy to understand: The model has never seen an input sequence in the Block-Attention manner during the training process. Therefore, the sharp drop makes sense. Next, we will focus on the second conclusion. From Table 1, we can easily find that the models trained with two attention methods, namely Llama3-vanilla-sft and Llama3-block-ft, not only have comparable effects on in-domain test sets but also have little difference on out-domain test sets. This indicates that in the RAG scenario, it is completely feasible to replace self-attention with Block-Attention, and there will be almost no performance loss.
Moreover, the experimental results on Mistral show some interesting results: the accuracy of Block-Attention models actually slightly outperformed the results of self-attention models on most datasets. It raised a conjecture: is Block-Attention a better attention mechanism than the self-attention mechanism? Intuitively, the semantics of each retrieved passage are mutually independent, and the self-attention method may cause the representation of the passage to be interfered by the context and not accurate enough. However, we still need sufficient experiments to verify the correctness of this assertion, so this is only a conjecture and not our conclusion.
@@
Finally, one may still be interested in knowing exactly how many training steps are needed for the model to adapt to the Block-Attention mechanism. Therefore, we counted the accuracy of the Llama3-8B model on the four test sets at different fine-tuning steps and plotted it in Figure 4. It can be observed that at the beginning stage of fine-tuning, there is a huge performance difference between the two models. It makes sense because the model needs more training steps to adapt to the Block-Attention manner. After about 400 training steps, the model completely adapts to the new attention mechanism, and the two models show comparable accuracy.
@@ 3.6 Inference Efficiency of Block-Attention
In the previous section, we already addressed our first concern: After fine-tuning, the Block-Attention model can achieve similar or even better performance than the self-attention model. In this section, we focus on our second concern: How much can the Block-Attention mechanism reduce the TTFT and FLOPs-TFT?
@@ 3.7 Discussion
Furthermore, some dynamic RAG paradigms such as ReAct (Yao et al., 2023) have been proven to enhance the general task-solving ability of the LLM. However, compared to vanilla RAG, such paradigms need to iteratively compute the KV states of retrieved passages for single user input, which results in a disastrous inference latency. With the help of Block-Attention, the additional latency consumption in this aspect can be reduced to a level that users can tolerate. Developers will be able to safely use such paradigms in online scenarios to optimize the reasoning ability of LLM.
@@ 4.1 Retrieval-Augmented Generation (RAG)
RAG is a widely used technique to improve generations of language models by using retrieved nearest-neighbor documents or passages as references, which typically involves two stages: retrieval and generation. Before generation, retrieval finds most similar passages with the user query or the context, by using BM25 or dense retrieval model (Lee et al., 2021; Lan et al., 2023; 2024). After collecting retrieved passages, there are numerous techniques to incorporate the knowledge during generation. Earlier works includes concatenation (Izacard & Grave, 2021) and cross-attention (Borgeaud et al., 2022). Some studies have also begun to attempt to directly use the retriever as text generator, that is, text generation is performed by selecting context-aware phrases from a collection of supporting documents.  (Lan et al., 2023; Cao et al., 2024)
@@
4.2 Efficiency of RAG
The efficieny of LLM inference gradually decreases as the sequence length of the concatenation of retrieved passages rapidly increases, leading to a huge challenges for retrieval augmented generation of LLMs. To solve this problem, there are numerous novel architectures, model quantization and inference strategies are proposed. For example, the linear attention and Mixture of Expert (Fedus et al., 2022) architecture are widely used architectures for reducing the complexity of LLM inference (Shazeer, 2019; Zhang & Cai, 2022; Fedus et al., 2022). Speculative decoding leverages the parallel of language models to speed up the inference (Leviathan et al., 2023). However, most previous works are not designed for RAG scenarios. Recently, RAG Cache(Jin et al., 2024) caches the intermediate states of retrieved knowledge in GPUs, showing 4x TTFT.
Another stream of researches focus on individually and parallely process each documents, which is related to our work, such as FiD (Izacard & Grave, 2021), PCW (Ratner et al., 2023a) and Prompt Cache (Gim et al., 2024). However, FiD is widely used for encode-decoder architecture, and not compatible for the decoder-only architecture of LLMs, since they need to concatenate the hidden states for decoder during inference. PCW focuses on extending the context window rather than efficient inference. Prompt Cache enables the efficient inference for RAG applications by designing the Prompt Markup Language (PML), while their pre-computed KV cache of retrieved passages are bound to the specific IDs, leading to its limited applications. In other words, once the rank of a passage in the input sequence changes, its KV cache cannot be reused.
@@ 5 Conclusion
We also want to note that Block-Attention plays an important role in many scenarios and is not limited to RAG only. We only introduced its effects on RAG because of that, due to some confidentiality reasons, we cannot disclose how we use it in our industrial applications. We look forward to researchers from the community being able to further explore the potential of Block-Attention and apply it to more appropriate scenarios.

### v5

@@ 3.5 Main Results
From the results in Table 2, we can draw four important conclusions:
1) It is not advisable to directly switch from full-attention to parallel prompt encoding methods including Block-attention, as it will lead to a sharp drop in accuracy. For instance, as can be seen from the experimental results of Tulu3-block-w/o-ft, removing the Block fine-tune process causes the Tulu3-RAG model to experience an average absolute performance decrease of 16.2% across all four RAG benchmarks. Additionally, the performance degradation of Tulu3-promptCache and Tulu3-Superposition is even more significant, which also indicates that the current parallel prompt encoding methods are still far from satisfactory. Through case studies, we found that these models have serious degeneration problems.
2) However, if we use the Block-attention mechanism in the fine-tuning stage, then the resultant model has comparable performance with its full-attention counterpart. In RAG scenarios, the performance gap between Tulu3-block-ft and Tulu3-RAG on four benchmarks is less than or equal to 1%. Meanwhile, Tulu3-block-ft significantly outperforms Tulu3-SFT on 2wiki and HQA. In the ICL scenario, Tulu3-block-ft also has an overall performance almost the same as or even slightly higher than that of the two full-attention models. This conclusion indicates that in RAG and ICL scenarios, it is completely feasible to replace full-attention with Block-attention, and there will be almost no performance loss.
3) Block and full attention can seamlessly transition. When we switch Tulu3-block-ft to the full-attention mode, we are pleasantly surprised to find that, whether in the RAG scenario (denoted as Tulu3-block-ft-full) or in the naturally single-block zero-shot scenario (IFEval, HumanEval, and MMLU), the Block-attention model can achieve performance comparable to or even slightly higher than that of the two strong full-attention baselines. This conclusion, together with the previous one, indicates that using both Block-attention and full-attention during training can lead to the acquisition of a model that is capable of seamlessly transitioning between the two attention mechanisms.
@@
4) The position re-encoding operations are essential for the Block-attention Model. Removing it leads to a certain degree of performance drop—an average 2% decrease in accuracy on all RAG datasets. Additionally, in this case, the model occasionally experiences the problem of degeneration.
Figure 4: The accuracy of model checkpoint
@@
Finally, one may still be interested in knowing exactly how many training steps are needed for the model to adapt to the Block-attention mechanism. Therefore, we counted the average accuracy of Tulu3-block-ft and Tulu3-block-ft-full on four RAG benchmarks at different fine-tuning steps and plotted it in Figure 4. It can be observed that at the beginning stage of fine-tuning, there is a huge performance difference between the two attention modes. It makes sense because the model needs more training steps to adapt to the Block-attention manner. After about 800 training steps, the model can switch seamlessly between the two attention modes without any performance loss.
@@ 3.6 Inference Efficiency of Block-attention
In the previous section, we already addressed our first concern: After fine-tuning, the Block-attention model can achieve similar or even better performance than the full-attention model. In this section, we focus on our third concern: How much can the Block-attention mechanism reduce the TTFT and FLOPs-TFT?
@@ 3.7 Discussion
While our experiments in this paper focus on the RAG and ICL scenario using publicly available datasets, Block-attention’s transformative potential extends far beyond these domains. Through validation across multiple internal tasks, we observed consistent efficiency gains. However, due to confidentiality constraints preventing data/code disclosure, we strategically selected the RAG scenario for reproducibility and benchmarking purposes.
Notably, Appendix A details how Block-attention addresses critical latency challenges in Game AI. We believe that Block-attention’s true disruptive potential lies in enabling real-time LLM agents - a vision technically unattainable through full-attention due to the constraints of inference costs.
@@ 4 Related Works
4.1 Concurrent Work
We were surprised to find that another ICLR 2025 submission, TurboRAG (Lu & Tang, 2025), independently proposed similar methods and reached similar conclusions as ours. They also proposed independent attention (which is the block attention in this paper) and reordered positions (the position re-encoding in this paper). Unfortunately, their work was not accepted by ICLR 2025. We sincerely hope that the community will also recognize them as one of the pioneers of Block-attention, and we hope that this similarity will not affect the publication of their paper in the future. Another study, DecoupledRAG (Dong et al., 2025), addresses the inefficiency issues in traditional Retrieval-Augmented Generation (RAG) methods by employing a cross-attention mechanism to directly inject external knowledge into the LLM’s reasoning process.
Recently, two highly discussed papers in the sparse attention domain—NSA (Native Sparse Attention) from Deepseek (Yuan et al., 2025) and MoBA (Mixture of Block Attention) from Moonshot (Lu et al., 2025)—have independently proposed block-based attention mechanisms similar to our Block-attention. While these works are not strictly parallel to ours (their preprints were released five months after ours), we aim to clarify the distinctive contributions and orthogonality of our approach compared to these studies. After partitioning inputs into blocks, these methods employ a trainable block selection operation to filter out irrelevant blocks. The retained blocks are then concatenated and processed in an auto-regressive pre-filling manner. Our Block-attention, in the contrary, focuses on two novel aspects: 1) Parallel context encoding, and 2) Cross-prompt block KV cache reuse. Neither NSA nor MoBA supports these capabilities.
In our prior work Zhang et al. (2024), we demonstrated that even simple block selection mechanisms—when applied to parallel context encoding—can substantially improve model accuracy. This suggests the “Block-attention we proposed” addresses complementary optimization dimensions compared to the “Block-attention proposed by DeepSeek and Moonshot”. In their future models, they may take our work as another pathway to significantly reduce their inference cost further.
@@ 4.2 Retrieval-Augmented Generation (RAG)
RAG is a widely used technique to improve generations of language models by using retrieved nearest-neighbor documents or passages as references, which typically involves two stages: retrieval and generation. Before generation, retrieval finds most similar passages with the user query or the context, by using BM25 or dense retrieval model (Lee et al., 2021; Lan et al., 2023; 2024; Ma et al., 2025; Che et al., 2024). After collecting retrieved passages, there are numerous techniques to incorporate the knowledge during generation. Earlier works includes concatenation (Izacard & Grave, 2021), cross-attention (Borgeaud et al., 2022) and distribution interpolation Khandelwal et al. (2020). Some studies have also begun to attempt to directly use the retriever as text generator, that is, text generation is performed by selecting context-aware phrases from a collection of supporting documents (Lan et al., 2023; Cao et al., 2024).
@@
4.3 Parallel Context Encoding
Some researches focus on individually and parallely process each documents, which is related to our work, such as SGLang (Zheng et al., 2024), FiD (Izacard & Grave, 2021), PCW (Ratner et al., 2023), PromptCache (Gim et al., 2024) and CacheBlend (Yao et al., 2024). Although SGLang could maintain the comparable generation quality as the original model, its reuse conditions are very strict, making it difficult to improve the inference efficiency. FiD is widely used for encode-decoder architecture, and not compatible for the decoder-only architecture of LLMs, since they need to concatenate the hidden states for decoder during inference. PCW focuses on extending the context window rather than efficient inference. The acceleration effect of PromptCache is similar to Block-Attention; however, as shown in our experiments, due to not properly handling positional encoding, its RAG performance is rather poor. CacheBlend introduces a trade-off between generation quality and KV cache reuse efficiency. Compared to SGLang (Zheng et al., 2024), CacheBlend achieves higher reuse efficiency and lower TTFT but at the cost of slightly reduced generation quality. On the other hand, compared to PromptCache, CacheBlend incurs overhead due to cross-attention recovery but delivers better generation quality.
From the above analysis, one may derive an intriguing conclusion: existing studies are forced to compromise either on generation quality or inference efficiency. Some works maintain high quality with low efficiency (SGLang), others prioritize high efficiency (PromptCache), while some attempt to balance both aspects (CacheBlend). In contrast, our proposed Block-Attention eliminates the need for such trade-offs through position re-encoding and block fine-tuning. It achieves both the same generation quality as the original model and the reuse efficiency comparable to PromptCache, without introducing additional overhead. More importantly, the resulting model even achieves seamless switching between Block-Attention and full-attention modes, significantly enhancing the flexibility of online services.
@@ 5 Conclusion
6 Acknowledgements
We sincerely express our gratitude to our friends, JCY, ZZL, HWY, and QXT, for their insightful suggestions and unwavering support during the inception of this idea. Unfortunately, due to their affiliation with a confidential institution, we are unable to include them in the author list. We earnestly hope that they will eventually enjoy a more open research environment.
We also extend our heartfelt thanks to our colleagues at Tencent AI Lab—Xinting Huang, Tian Liang, Jiahao Xu, and Zhaopeng Tu—for their valuable advice and resource support during both the rebuttal stage and the camera-ready phase of this work.
Appendix A Block-attention in Game AI
Before detailing the application of Block-attention in Game AI, we first outline the characteristics of LLM tasks in gaming scenarios. In such contexts, the LLM input typically consists of the game’s current state, known as gamecore data in the industry. This data is structured as JSON with lengths ranging from thousands to hundreds of thousands of tokens. Different tasks produce varied outputs, but the input generally remains the gamecore data of the current frame. For example: for AI players like AlphaStar (Vinyals et al., 2019) and JueWu (Ye et al., 2020), their problem can be formulated as:
$P(Player\>Action|gamecore).$
For AI NPCs, their problem then becomes:
$P(NPC\>Action|gamecore).$
As for tasks like AI commentary, they can be defined as:
$P(Game\>Comments|gamecore).$
As shown in the formulations above, most gaming tasks involve extremely long contexts (from several K to hundreds of K) but relatively short output sequences (tens to hundreds of tokens). If the AI must prefill data from scratch for every decision, the resulting excessive inference latency would render LLMs impractical for real-time applications. This explains why current MOBA and FPS games still rely on small-parameter models for online deployment.
Figure 5: A case of Texas hold’em AI’s gamecore data
To illustrate how Block-attention reduces TTFT in LLMs, consider a Texas Hold’em example. Figure  5 compares gamecore data at two consecutive states: State N-1 (previous) and State N (current). Notably, these states are nearly identical except for the value of state[’chips’][’p2’], which changes from {’bet’: 10, ’remain’: 990} to {’bet’: 50, ’remain’: 950}. This high inter-frame repetition (exceeding 99.5% based on our analysis) is common in games. Furthermore, the structured JSON format of gamecore data allows rule-based partitioning into independent blocks. By encoding only new or modified blocks during prefilling (instead of prefilling from scratch), we achieve equivalent accuracy to full prefilling while significantly reducing latency.
In an unreleased game (not yet available to the public), by dividing the game state into over 300 independent blocks, we achieved the same accuracy as full-attention mechanisms in these tasks, while cutting the average TTFT from 2,800ms to 100ms and reducing query latency from 3,000ms to under 300ms. Our Block-attention method aims to help AI game developers effortlessly create more games powered by advanced AI systems. We envision that Block-attention will empower AI researchers in the gaming industry to seamlessly develop more AI-centric gaming experiences.